# Ctrl-Room: Controllable Text-to-3D Room Meshes Generation with Layout Constraints

### v4

@@ 9 Mask-Guided Editing
For conciseness, we denote the target image $\hat{\mathbf{x}}^{\rm edit}_{0}$ initialized by $\hat{\mathbf{x}}^{\rm new}_{0}$. We first propagate the original image $\mathbf{x_{0}^{\rm ori}}$ and $\hat{\mathbf{x}}^{\rm edit}_{0}$ to get $\mathbf{x_{t}^{\rm ori}}$ and $\hat{\mathbf{x}}^{\rm edit}_{t}$ at timestep $t$ by diffusion process, respectively. At each iteration, we use the same ControlNet model to denoise both $\mathbf{x_{t}^{\rm ori}}$ and $\hat{\mathbf{x}}^{\rm edit}_{t}$ and extract the latent features of them, denoted as $F^{\rm ori}_{t}$ and $F^{\rm edit}_{t}$, respectively. Based on the strong correspondence between the features, the source mask area $\mathbf{m}_{\rm src}$ and the target area $\mathbf{m}_{\rm tar}$ in $F^{\rm ori}_{t}$ and $F^{\rm edit}_{t}$ need to have high similarity. Here, we utilize the cosine embedding loss to measure the similarity, and define the optimization loss function as follows:
$\mathcal{L_{\rm obj}}=-cos({\rm sg}(F^{ori}_{t}\odot\mathbf{m}_{\rm src}),F^{edit}_{t}\odot\mathbf{m}_{\rm tar})$ (17)
Here, ${\rm sg}$ is the stop gradient operator, the gradient will not be back-propagated for the term ${\rm sg}(F^{ori}_{t}\odot\mathbf{m}_{\rm src})$. Then we minimize the loss iteratively. At each iteration, $\hat{\mathbf{x}}^{\rm edit}_{t}$ is updated by taking one gradient descent step with a learning rate $\eta$ to minimize the loss $\mathcal{L_{\rm obj}}$ as,
@@ 11 Implementation details
  • In the layout generation stage, We train the scene code diffusion model on our processed typical indoor rooms data of Structured3D [48] for $200,000$ steps. The frozen text encoder we adopted is the same as Stable Diffusion [24]. The training is performed using the AdamW optimizer with a batch size of $128$ and a learning rate of $41e-$ 4, utilizing $2$ A6000 GPUs. During the inference process, we utilize the DDIM [31] sampler with a step size of $200$ to perform scene code denoising.
@@
  • As for the mask-guided editing module, we utilize the fine-tuned Control-Seg model to inpaint the background content and optimize the latents of the edited panorama. In inpainting step, the weights used too fuse the unpainted area and unchanged area are set $\lambda_{\rm ori}=0.8,\lambda_{\rm new}=0.2$ In the optimization step, the maximum iteration is $M=50$, the learning rate $\eta$ for optimization is initialized to $0.1$ and then gradually decreases to $0.01$.
@@ 11.1 Baseline Implementations
  • MVDiffusion [36]: To get a high-resolution photo realistic panorama, MVDiffusion employs $8$ branches of SD [24] model and correspondence-aware attention mechanism to generate multi-view images simultaneously. We first fine-tune the pre-trained model of MVDiffusion on Structured3D for 10 epochs(about 3 days). Since each generated subview image of MVDiffusion is at $512\times 512$ resolution, the final panorama is pretty large. We resize the generated panorama of MVDiffusion from $4096\times 2048$ to $1024\times 512$. Then the 8 subview perspective images are extracted from the post-processed panorama using the same camera settings (FOV=$90^{\circ}$,rotation=$45^{\circ}$). The same operation is adopted on our generated panoramic images. Finally, we combine the panorama from MVDiffusion with the depth estimation [45] and Poisson reconstruction [14] module to create a 3D mesh.
@@ 14 User Study
Follow Text2Room [12], we conduct a user study and ask $n=61$ ordinary users to score the Perceptual Quality(PQ) and 3D Structure Completeness(3DS) of the generated room on a scale of $1-5$. Different from Text2room which only demonstrates the perspective renderings of the 3D room, we directly show users the generated mesh to get a global evaluation of the whole generated 3D room. We show an example of the presented interface of the user study in Fig. 21. In total, we presented $40$ top-down views from $10$ scenes and report averaged results for each method. Users favor our approach, which emphasizes the superiority of our more plausible geometry, along with the vivid texture.

### v5

@@ 9 Mask-Guided Editing
For conciseness, we denote the target image $\hat{\mathbf{x}}^{\rm edit}_{0}$ initialized by $\hat{\mathbf{x}}^{\rm new}_{0}$. We first propagate the original image $\mathbf{x_{0}^{\rm ori}}$ and $\hat{\mathbf{x}}^{\rm edit}_{0}$ to get $\mathbf{x_{t}^{\rm ori}}$ and $\hat{\mathbf{x}}^{\rm edit}_{t}$ at timestep $t$ by diffusion process, respectively. At each iteration, we use the same ControlNet model to denoise both $\mathbf{x_{t}^{\rm ori}}$ and $\hat{\mathbf{x}}^{\rm edit}_{t}$ and extract the latent features of them, denoted as $F^{\rm ori}_{t}$ and $F^{\rm edit}_{t}$, respectively. Based on the strong correspondence between the features, the source mask area $\mathbf{m}_{\rm src}$ and the target area $\mathbf{m}_{\rm tar}$ in $F^{\rm ori}_{t}$ and $F^{\rm edit}_{t}$ need to have high similarity. Here, we utilize the cosine embedding loss to measure the similarity, and define the optimization loss function as follows: 6. Here, $\rm sg$} is the stop gradient operator, the gradient will not be back-propagated for the term ${\rm sg}(F^{ori}_{t}\odot\mathbf{m}_{\rm src})$. Then we minimize the loss iteratively. At each iteration, $\hat{\mathbf{x}}^{\rm edit}_{t}$ is updated by taking one gradient descent step with a learning rate $\eta$ to minimize the loss $\mathcal{L_{\rm obj}}$ as,
@@ 11 Implementation details
  • In the layout generation stage, We train the scene code diffusion model on our processed typical indoor rooms data of Structured3D [48] for $200,000$ steps. The frozen text encoder we adopted is the same as Stable Diffusion [24]. The training is performed using the AdamW optimizer with a batch size of $128$ and a learning rate of 1e41e-4, utilizing $2$ A6000 GPUs. During the inference process, we utilize the DDIM [31] sampler with a step size of $200$ to perform scene code denoising.
@@
  • As for the mask-guided editing module, we utilize the fine-tuned Control-Seg model to inpaint the background content and optimize the latents of the edited panorama. In inpainting step, the weights used too fuse the unpainted area and unchanged area are set $\lambda_{\rm ori}=0.8,\lambda_{\rm new}=0.2$ . In the optimization step, the maximum iteration is $M=50$, the learning rate $\eta$ for optimization is initialized to $0.1$ and then gradually decreases to $0.01$.
@@ 11.1 Baseline Implementations
[…]
@@ 14 User Study
Follow Text2Room [12], we conduct a user study and ask $n=61$ ordinary users to score the Perceptual Quality(PQ) and 3D Structure Completeness(3DS) of the generated room on a scale of 151-5. Different from Text2room which only demonstrates the perspective renderings of the 3D room, we directly show users the generated mesh to get a global evaluation of the whole generated 3D room. We show an example of the presented interface of the user study in Fig. 21. In total, we presented $40$ top-down views from $10$ scenes and report averaged results for each method. Users favor our approach, which emphasizes the superiority of our more plausible geometry, along with the vivid texture.